### v1

# citations.bib

Abstract
Detecting vulnerabilities within compiled binaries is challenging due to lost high-level code structures and other factors such as architectural dependencies, compilers, and optimization options. To address these obstacles, this research explores vulnerability detection by using natural language processing (NLP) embedding techniques with word2vec, BERT, and RoBERTa to learn semantics from intermediate representation (LLVM) code. Long short-term memory (LSTM) neural networks were trained on embeddings from encoders created using approximately 118k LLVM functions from the Juliet dataset. This study is pioneering in its comparison of word2vec models with multiple bidirectional transformer (BERT, RoBERTa) embeddings built using LLVM code to train neural networks to detect vulnerabilities in compiled binaries. word2vec Continuous Bag of Words (CBOW) models achieved 92.3% validation accuracy in detecting vulnerabilities, outperforming word2vec Skip-Gram, BERT, and RoBERTa. This suggests that complex contextual NLP embeddings may not provide advantages over simpler word2vec models for this task when a limited number (e.g. 118K) of data samples are used to train the bidirectional transformer-based models. The comparative results provide novel insights into selecting optimal embeddings for learning compiler-independent semantic code representations to advance machine learning detection of vulnerabilities in compiled binaries.
Index Terms:
Machine Learning, Buffer Overflows, BERT, RoBERTa, Binary Security, LLVM, word2vec
I Introduction
Microsoft is one of the dominant players within the server and desktop operating system market; therefore, vulnerabilities in its software can negatively impact millions of systems around the world. Some of the most prolific worms in the world have exploited vulnerabilities in software developed by Microsoft; these worms include Code Red [RN72], which exploited a vulnerability in Microsoft IIS, Sasser [RN73], which exploited the Windows LSASS process, and Conficker [RN74] and WannaCry [RN75] which exploited SMB vulnerabilities in Windows systems. The financial impact of recovering from these worms is estimated to be billions of dollars (USD) [RN76]. Microsoft’s closed-source nature  [RN77] prevents organizations from reviewing code written in a high-level programming language for vulnerabilities before deploying it within their environments. Microsoft is a single example of an organization that produces closed-source products, but is certainly not the only company. Most commercial organizations protect their intellectual property by not releasing their source code to the public. Therefore, organizations that use these commercial products rely entirely on the organizations that create the software to follow secure coding practices. However, they cannot independently verify the security of the code.
Locating vulnerabilities within compiled binaries can be difficult, because significant information is lost when source code is compiled on a target platform. The names of variables, programming structures, and code comments are reduced to low-level machine code. Furthermore, machine code is significantly different for each platform on which it is compiled (e.g., x86/x64, MIPS, SPARC, ARM, etc.). To solve these issues, many researchers have turned to machine learning in the hope that it may provide a solution for locating vulnerabilities in compiled binaries.
Due to perceived similarities between natural languages and code, researchers investigating machine learning for vulnerability detection in compiled binaries have investigated using natural language processing (NLP) techniques for this purpose. The current body of knowledge extensively utilized neural networks that are commonly used within NLP, including LSTM, gated recurrent unit (GRU), bidirectional LSTM (BLSTM), and bidirectional GRU (BGRU) neural networks. Furthermore, NLP encoding models such as word2vec, instruction2vec, BERT, and RoBERTa were used to learn semantic relationships within the code. The embeddings these models generate are commonly used to train neural networks to detect vulnerabilities in decompiled code. A summary of this previous work is given in Section  V.
Although researchers have explored the use of specific natural language processing (NLP) encoding techniques in the context of decompiled code, research has not yet focused on comparative analysis using the most popular NLP encoding models to identify which embedding techniques perform best in the task of identifying vulnerabilities in compiled binaries. Very few studies have used bidirectional transformer-based models to acquire semantics within decompiled code for this task. RoBERTa was used by [RN67] to learn semantics within the x86 assembly language, and [gallagherllvm] used this model to learn the semantics of the LLVM code. Additionally, [koo2021semantic] uses BERT for code similarity detection, and [han2022binary] trained a BERT model using pcode and used the embeddings to differentiate between vulnerable and non-vulnerable code. Besides [gallagherllvm] and [han2022binary], little to no research has explored the use of unidirectional encoders to identify vulnerabilities in compiled binaries that have been lifted to languages such as LLVM. Furthermore, according to our research, there have been no comparative studies on the performance of neural networks trained using embeddings from common NLP models (word2vec, BERT, and RoBERTa) to differentiate between vulnerable and non-vulnerable LLVM code. This paper explores the use of multiple bidirectional encoding models (BERT & RoBERTa) to evaluate their effectiveness in generating embeddings that can be used to accurately detect vulnerabilities in compiled binaries. The models built using these bidirectional embedding models are compared to those generated using word2vec: CBOW and Skip-Gram embeddings.
This paper can be viewed as an extension of previous work of [RN9] and [schaad2022deeplearningbased]. In these studies, the researchers compiled code samples from the SARD dataset, lifted them to LLVM using the RetDec tool, and used the lifted code to build a word2vec model. Once this model was built, the embeddings from the word2vec model were provided to neural networks to train them to differentiate between vulnerable and non-vulnerable code. However, the research articulated in this paper acknowledges the limitations of word2vec’s ability to create context-aware embeddings and explores the alternative use of bidirectional transformer-based models.
The remainder of the paper is organized as follows. Section II details the design and implementation, and section III presents the results. Section V describes the work related to the current research. Section VI presents ideas for future work, followed by the conclusion in Section VII.
II Design and Implementation
The current research uses machine learning techniques to identify vulnerabilities in compiled binaries by first using NLP bidirectional encoders to learn code-level semantics in lifted LLVM code. The embeddings generated from these encoders were then used to train LSTM neural networks, and the performance of these neural networks was compared with neural networks trained using simpler word2vec embedding models. An overview of the steps involved in this process is as follows.
  1. 1.
    All code samples within the NIST SARD [RN11] Juliet dataset were compiled into object files, and the object files were lifted to LLVM. Although the Juliet dataset contains several classes of vulnerabilities and a mix of code samples containing vulnerable and non-vulnerable functions, the primary purpose of the dataset was to learn the semantics between tokenized LLVM instructions. Therefore, tracking vulnerability classes and vulnerable/non-vulnerable status was not required. This process is located in II-A.
  2. 2.
    A second dataset was created that contained vulnerable and non-vulnerable LLVM functions generated using samples containing Stack-Based Buffer Overflow (CWE-121) vulnerabilities taken from the Juliet dataset. The vulnerability status of each of these code samples was tracked within this dataset. The process of creating this data set is described in II-B.
  3. 3.
    The LLVM functions that comprised both datasets were isolated and processed to convert the code into tokens that resemble a natural language. The goal of this phase was to convert the raw LLVM code to a format that could be easily provided to an NLP embedding model. The preprocessing performed is described in II-C.
  4. 4.
    The processed LLVM code in the first dataset was used to train four NLP embedding models. Two of these embedding models, BERT and RoBERTa, are well-known NLP bidirectional models, and two are much simpler embedding models used extensively throughout the current literature. The creation of the word2vec model is covered in II-D, and the creation of bidirectional encoders is covered in II-E.
    • BERT
    • RoBERTa
    • word2vec: CBOW embedding
    • word2vec: Skip-Gram embedding
  5. 5.
    The second dataset that comprised both vulnerable and non-vulnerable functions was provided to each embedding model to generate embeddings sent to the LSTM neural network to train the model to detect functions containing CWE-121 vulnerabilities. The LSTM creation process is described in II-F.
  6. 6.
    Once each LSTM network was trained, the performance of the model was evaluated using accuracy, precision, recall, and the F1 score. The performance of each LSTM is described in III.
II-A Embedding Dataset Creation
The dataset chosen for this study is the C/C++ code samples from the Juliet 1.3 dataset [RN89] within the larger SARD dataset [RN11]. The rationale for selecting this dataset lies in its use in several other research projects [RN50, RN43, RN67, RN56, RN47, RN46, RN55, schaad2022deeplearningbased]. Choosing a dataset used within multiple research projects provides a basis for a meaningful comparison with the previous corpus of knowledge, as well as for easier reproducibility of results.
The Juliet dataset is comprised of source code samples for several classes of vulnerabilities. For each vulnerability, the dataset contains both a vulnerable and a non-vulnerable version of the source code, and can thus be used to train machine learning models to detect various vulnerabilities.
The function names within the code make it easy to differentiate between the vulnerable and non-vulnerable versions of the code. The non-vulnerable versions of each function have the word Good within them, and the vulnerable versions of the code have the word Bad in them. Furthermore, these functions may call secondary functions, such as GoodSink or GoodSource.
The Juliet data set, which consists of 101,231 C/C++ files, was compiled to separate object files on a Kali 2023.3 and Windows 10 system. Once each source code sample was compiled into an object file, the RetDec tool was used to convert the object files into an LLVM file. Many articles analyzed during the literature review described research on building models to detect vulnerabilities within x86/x64 assembly language instructions. However, this approach limits the applicability of the results to that instruction set, and the same process cannot be applied to additional architectures. Thus, the current research project used RetDec [RN22] to lift binaries from x86/x64 to an intermediary language called LLVM. The compiler-agnostic [RN18] nature of LLVM allows for future expansion of datasets for samples compiled using different architectures. Additionally, the use of LLVM should help reduce some differences between compiler-specific optimizations [RN18]. Theoretically, a lifted binary for x86/x64 should be identical to a lifted binary compiled for a different architecture (e.g., ARM, SPARK, etc.).
This first dataset included all functions from the lifted binaries and served the purpose of training embedding models to grasp the semantics of the code level. The breakdown of the number of functions within this dataset is given in Table I.
TABLE I: Embedding Training Dataset
Operating System Object Files LLVM Files
Extracted
Functions
Windows 10 96,061 94,225
Kali Linux
2023.3
65,972 65,972
Total
(Pre Duplicate
Removal)
162,033 160,197 2,479,856
Total
(Post Duplicate
Removal)
117,742
II-B Neural Network Dataset Creation
Creating the second dataset involved isolating the Stack-Based Buffer Overflow (CWE-121) LLVM samples from the dataset used to train the embedding models. The breakdown of the number of functions within this dataset is in Table  II.
TABLE II: LSTM Training Dataset
Operating System Object Files LLVM Files
Extracted
Functions
Windows 10 7,871 7,009
Kali Linux
2023.3
4,949 4,949
Total
(Pre Duplicate
Removal)
12,820 11,958
31,301
-17,501 (Clean)
-13,800 (Vuln)
Total
(Post Duplicate
Removal)
6,888
-4,319 (Clean)
-2,569 (Vuln)
Total
(Removal of
Functions More
than 2048 Tokens)
6,773
-4,227 (Clean)
-2,546 (Vuln)
Creating the dataset of vulnerable and non-vulnerable functions presented notable challenges due to the naming convention employed by the Juliet dataset. In some cases, the base function (e.g., FunctionName_Bad) may have a vulnerability embedded within the function itself. However, in other cases, the function representing the vulnerable or non-vulnerable version of the code may be a simple function that calls secondary functions that contain the vulnerability (e.g., Function_Name_BadSink). Additionally, in many cases, the non-vulnerable version of the code contained a function that included the keyword Good2Bad or Bad2Good. Thus, when building a dataset representing vulnerable and non-vulnerable versions of the code, all functions containing keywords used to indicate a non-vulnerable version of the code (i.e., Good, goodG2B, goodB2G) and vulnerable versions of the code (i.e., Bad) were included in the dataset. To reduce the number of functions with the sole purpose of calling secondary functions, duplicate versions of each function were removed after the preprocessing articulated in section  II-C.
Tokens from both vulnerable and non-vulnerable functions were fed into NLP embedding models to create embeddings for training LSTM models. Bidirectional encoding models like BERT and RoBERTa require a max_length parameter, restricting the number of tokens per processed function. Due to constraints in GPU memory, setting a max_length significantly beyond 2048 was not feasible. Consequently, samples exceeding the specified max_length were excluded from the dataset during the training and validation of neural networks.
II-C LLVM Preprocessing
Once the LLVM datasets were generated, a sampling of the LLVM files was manually analyzed to determine the best processing approach. The primary goal of the embedding algorithms was to learn the semantics of the LLVM code. Thus, preprocessing was performed to reduce the number of tokens to focus more on execution order and less on function-specific code offsets. A Jupyter Notebook was created that performs the actions listed in Table  III.
TABLE III: Preprocessing Steps
Action Before After
LLVM functions
created within the
code were given a
unique identifier
@_func_unique_name() func_x ()
global variables,
labels, and stack
pointers were renamed
to generic tokens
@global_var_offset
dec_label_pc_offset:
%stack_var_offset
globalvar_x
declabel_x
stackvar_x
Variables surrounded
by quotes were
converted to unique
tokens used to track
order of appearance
@“misc_var” “stringvar_x”
Numeric constants
greater than nine
were given a unique
identifier based on
order of appearance
12345
-12345
instconstpos_x
instconstneg_x
Special characters
were renamed to text-
based tokens
opencurlybrackets
dotpuncmark
commapuncmark
Note: x is a number used to track the order of appearance in the function.
II-D word2vec: CBOW and Skip-Gram Model Creation
As previously stated, the research team used word2vec: CBOW and Skip-Gram embeddings, BERT, and RoBERTa for generating embeddings. word2vec was selected due to its prevalence in the current body of research [RN68, RN57, RN36, RN7, RN9, RN58, schaad2022deeplearningbased], and its ability to provide a strong comparison point with the more complex NLP bidirectional encoders. The techniques used by word2vec – CBOW and Skip-Gram – and the code that implements them were released in 2013 [mikolov2013efficient].
The first of these techniques, CBOW, generates a vector that uniquely represents a particular word using the words before and after. It is especially good at identifying similarities between words because similar words are often found within the same word patterns. For example, in the sentence “one of my favorite foods to eat is ice cream.”, the word ice could easily be replaced with whipped or even sour. Thus, these words would have vectors that highlight that they are similar. However, the word road would have a vector that is quite different than ice, whipped, or sour because the words before and after the word are generally quite different from those for the other words. The CBOW technique predicts a missing word using the words surrounding the target word. The second technique used by word2vec, Skip-Gram, attempts to predict surrounding words using a target word, unlike CBOW, which predicts a target word based on the surrounding words.
The word2vec: CBOW and word2vec: Skip-Gram models were created using the gensim Python library. Each preprocessed LLVM function was tokenized using the word tokenizer function of the nltk library. The model was built using all the 117,742 samples from the embedding model dataset. There were 5,224 unique tokens used to train these models, but the vector size of the output embedding was set to one hundred. The window size was set to five and the minimum count required for each token was set to one.
II-E BERT and RoBERTa Model Creation
The code for the BERT model was released in 2019 [devlin2019bert] and stands for Bidirectional Encoder Representations from Transformers. BERT is a bidirectional encoder that learns token embeddings by masking out tokens within the input data and attempting to predict the missing token. Whereas some of the current models at the time were unidirectional, BERT proposed a bidirectional model that scored substantially higher on various NLP tasks than its predecessors. RoBERTa (Robustly Optimized BERT pre-training Approach) was also released in 2019 [liu2019roberta]. As its name suggests, RoBERTa is similar to BERT, with some optimizations used to increase its performance.
The custom BERT model used in the current research for learning semantics between LLVM instructions was created using the HuggingFace BertForMaskedLM library, and the BertWordPieceTokenizer was used to convert LLVM functions to tokens. Similarly, the RoBERTa model was generated using the RobertaForMaskedLM library and the ByteLevelBPETokenizer library. In both cases, 12 hidden layers with 100 nodes were used, each with an output layer of 100. The maximum vocabulary length was 2048, and batch sizes were limited to one due to GPU memory limitations. Training the BERT and RoBERTa models requires a training and validation dataset. Thus, the 117,742 samples from the embedding model dataset were split into two datasets; the training dataset comprised 90% of the overall dataset (i.e., 105,967) and the validation dataset was 10% of the dataset (i.e., 11,775).
For both the BERT and RoBERTa embedding models, the loss of training and validation steps was shown every 1,000 steps. The best loss in the validation set for the BERT model was achieved after step 42,000, and the best loss for the RoBERTa model was achieved at step 41,000 as shown in Tables  IV and  V. These were the embedding models used to create the embeddings provided to the LSTM models.
TABLE IV: BERT Five Best Loss Scores by Validation Step
Step Training Set Validation Step
42,000 2.698300 2.639552
41,000 2.694500 2.639749
40,000 2.689700 2.642417
38,000 2.701500 2.646453
39,000 2.690200 2.647793
TABLE V: RoBERTa Five Best Loss Scores by Validation Step
Step Training Set Validation Step
41,000 2.733500 2.701580
42,000 2.741900 2.703620
39,000 2.736500 2.708259
40,000 2.737100 2.709357
37,000 2.763900 2.712307
II-F LSTM Neural Network Configuration
This research uses embeddings generated from each encoding model to train LSTM neural networks to identify vulnerabilities within the code compiled. LSTM neural networks are recurrent neural networks designed to avoid vanishing gradients by using gates to track relevant information and forget irrelevant details. One of the advantages of recurrent neural networks is that they take the output from one step and reuse the output in subsequent steps. For example, suppose that an LSTM tries to predict the next word in the sequence, “my favorite food is ice…” There is a high probability that the next word is cream, so the LSTM would likely predict cream as the next word. However, if the next word entered is cold; the sentence has changed to “my favorite food is ice cold…” The LSTM can use the output of the previous step to update its prediction from cream to a different word such as custard. The recurrent nature of the LSTM should make it better suited for learning the patterns within the embeddings needed to identify vulnerable and non-vulnerable code. Although more complex neural networks may result in better performance, an LSTM was chosen because it is a good baseline neural network to evaluate the performance of different embedding models.
For the current research project, the LSTM neural networks comprised two hidden layers with 128 neurons each and 50 epochs to train each network. The Leaky Rectified Linear Unit (LeakyReLU) activation function was chosen because recurrent neural networks are prone to the problem of vanishing or exploding gradients [RN25, p. 335], and the LeakyReLU activation function helps avoid this problem. Each layer of the LSTM model was created using the TensorFlow and Keras libraries. A dropout rate of 20% was used between the LSTM layers and a single dense layer of a single neuron was used in the output layer. Furthermore, the layers of each embedding model were frozen so they could not be updated as the LSTM was being trained to differentiate between vulnerable and non-vulnerable code. Five LSTMs were generated for each embedding model. The first LSTM used the Adam optimizer with an initial learning rate of 0.0001. The second LSTM used the Stochastic Gradient Descent (SGD) optimizer with an initial learning rate and momentum of 0.01. Finally, the remaining three LSTM neural networks used the SGD optimizer with an initial learning rate of 0.0001 and momentums of 0.01, 0.001, and 0.0001. The dataset was randomized and divided into two subsets. The first subset was used to train each neural network and consisted of 80% (i.e., 5,418 samples) of the overall dataset; the second subset was used for model validation and consisted of 20% (i.e., 1,355) of the total dataset. The same training and validation subsets were used for each neural network.
III Results
Accuracy, precision, recall, and F1 score were evaluated for each LSTM neural network created from the embeddings generated from each embedding model.
III-A word2vec: CBOW
The metrics in Table  VI were recorded when each model achieved its highest accuracy. Although the highest accuracy, 92.3%, was obtained when the Adam optimizer was used with a learning rate of 0.0001, the model built using the SGD optimizer model with a learning rate and momentum of 0.01 was a close second with an accuracy of 91.7%. LSTMs created using the SGD optimizer with learning rates of 0.0001 had gradients that tended to stall in the middle epochs but started to spike in accuracy toward the latter epochs. However, in these final epochs, the model’s accuracy improved rapidly until the final epoch. Thus, it is likely that more epochs would have resulted in greater accuracy.
TABLE VI: word2vec: CBOW Metrics
Hyper-
parameters
Epoch Loss Accuracy Precision Recall
F1-
Score
Adam
-LR: 0.0001
50 0.2216 92.3% 87.9% 91.9% 89.9%
SGD
-LR: 0.01
-Mom: 0.01
49 0.2129 91.7% 88.2% 89.8% 89.0%
SGD
-LR: 0.0001
-Mom: 0.0001
50 0.4490 77.1% 66.5% 78.1% 71.9%
SGD
-LR: 0.0001
-Mom: 0.001
50 0.5164 72.2% 65.5% 54.5% 59.5%
SGD
-LR: 0.0001
-Mom: 0.01
50 0.5289 71.0% 65.1% 47.8% 55.2%
III-B word2vec: Skip-Gram
The metrics in Table  VII were calculated when the accuracy of the validation set was the highest for each model. The results for the SGD optimizer with a learning rate of 0.0001 and a momentum of 0.01, 0.001, and 0.0001 have been excluded because the model struggled to achieve accuracy gains beyond the base accuracy of placing all samples in a single category. The highest performing model built using word2vec Skip-Gram embeddings attained an accuracy of 90.6%.
TABLE VII: word2vec: Skip-Gram Metrics
Hyper-
parameters
Epoch Loss Accuracy Precision Recall
F1-
Score
Adam
-LR: 0.0001
40 0.2201 90.6% 87.4% 87.6% 87.5%
SGD
-LR: 0.01
-Mom: 0.01
30 0.3143 83.8% 74.9% 85.6% 79.9%
III-C BERT
The LSTM model that performed best with BERT embeddings was built using the Adam optimizer with a learning rate of 0.0001. Table  VIII highlights that the LSTM models trained using BERT embeddings hovered around an accuracy of 81%, with the best-performing model achieving an accuracy of 84.1%.
TABLE VIII: BERT Metrics
Hyper-
parameters
Epoch Loss Accuracy Precision Recall
F1-
Score
Adam
-LR: 0.0001
35 0.2985 84.1% 80.2% 76.4% 78.2%
SGD
-LR: 0.0001
-Mom: 0.001
42 0.3828 83.0% 74.3% 83.7% 78.7%
SGD
-LR: 0.01
-Mom: 0.01
16 0.3713 81.8% 70.0% 89.8% 78.7%
SGD
-LR: 0.0001
-Mom: 0.01
37 0.4003 81.3% 72.4% 81.1% 76.5%
SGD
-LR: 0.0001
-Mom: 0.0001
48 0.3975 81.0% 72.6% 78.9% 75.7%
III-D RoBERTa
Similarly to the LSTM models built using the BERT embeddings, the LSTM model with the best performance was achieved using the Adam optimizer with an initial learning rate of 0.0001. Table  IX highlights that neural networks created using RoBERTa embeddings consistently achieved better performance than those trained using BERT embeddings, with the highest performing model achieving an accuracy of 87%.
TABLE IX: RoBERTa Metrics
Hyper-
parameters
Epoch Loss Accuracy Precision Recall
F1-
Score
Adam
-LR: 0.0001
46 0.3152 87.0% 78.8% 89.4% 83.8%
SGD
-LR: 0.0001
-Mom: 0.01
45 0.3385 84.2% 73.9% 89.4% 80.9%
SGD
-LR: 0.0001
-Mom: 0.001
32 0.3737 83.3% 73.1% 87.6% 79.7%
SGD
-LR: 0.01
-Mom: 0.01
29 0.3489 83.0% 73.1% 86.6% 79.3%
SGD
-LR: 0.0001
-Mom: 0.0001
39 0.3721 82.9% 72.8% 86.8% 79.2%
IV Discussion
Neural networks created with the word2vec: CBOW embeddings achieved the best performance with a peak accuracy of 92.3% when the Adam optimizer was used with a learning rate of 0.0001. Similarly, models utilizing word2vec: Skip-Gram embeddings also demonstrated strong performance, achieving an accuracy of 90.6% with the same optimizer and learning rate. However, neither word2vec: CBOW and Skip-Gram performed well when the SGD optimizer was used with a lower learning rate (0.0001). Nevertheless, the behavior displayed by the word2vec CBOW accuracy and loss graphs indicate that additional epochs may have resulted in better performance.
The higher performance of word2vec CBOW over Skip-Gram is somewhat interesting. Once preprocessing was completed, the number of unique tokens within the LLVM functions was somewhat limited. This was due to the way the preprocessing limited unnecessary uniqueness within tokenized LLVM functions. The CBOW model’s technique tends to produce more accurate embeddings for representing tokens with higher frequency than the Skip-Gram model [mikolov2013efficient]. However, Skip-Gram models typically produce more accurate embeddings than CBOW models when trained with smaller datasets [mikolov2013efficient]. Thus, in this research, the accuracy gains achieved through token frequency resulted in better CBOW performance despite the relatively small training dataset.
The top accuracy of neural networks trained using BERT and RoBERTa embeddings were not as high as those trained using word2vec embeddings. This implies that, for smaller datasets, the simpler techniques used by word2vec outperform the more complex context-aware techniques used by BERT and RoBERTa. This result is not overly surprising considering the BERT model was pretrained using the BooksCorpus (800M words) and English Wikipedia (2,500M words) [devlin2019bert] and RoBERTa was trained on ten times that amount of data [liu2019roberta]. Thus, it is suspected that neural networks built using embeddings from BERT and RoBERTa models trained on larger datasets would likely achieve higher performance. However, despite the smaller training dataset used to train the BERT and RoBERTa models, it is interesting to note that neural networks utilizing these embeddings, optimized with SGD and a learning rate of 0.0001, outperformed the word2vec models trained under the same conditions. Furthermore, these models achieved a relatively strong top performance, achieving high accuracy rates of 87.0% with RoBERTa and 84.1% with BERT when the Adam optimizer was used with a learning rate of 0.0001. Thus, even though these models are meant to be trained with large amounts of data, they still manage to perform moderately well with smaller datasets.
V Related Work
The datasets integrated into the National Institute of Science and Technology’s (NIST) Software Assurance Reference Dataset [RN11] are among the most frequently referenced source code repositories in the reviewed literature. [RN9, RN10, RN57, RN45] explicitly stated that they used the NIST SARD dataset when selecting samples used for their research. One of the SARD subdatasets, the Juliet dataset, was used by [RN50, RN43, RN67, RN56, RN47, RN46, RN55, schaad2022deeplearningbased]. Two additional subdatasets within SARD used by researchers include MIT Lincoln Laboratories (MLL05b), used by [RN41], and STONESOUP (IARPA12, IARPA14), used by [RN43].
Training auxiliary models for learning semantical relationships in code is commonly used throughout the literature. [RN68, RN57, RN36, RN7, RN9, RN58, schaad2022deeplearningbased] used word2vec to generate embedding models to capture semantics within the code. However, due to the limitations of relying on word2vec to represent relationships within decompiled code, other studies propose alternative embedding models similar to word2vec or built upon word2vec embeddings. These alternatives include Instruction2vec [RN43], Bin2vec [RN47], VDGraph2Vec [RN67], BinVulDet [RN46], and VulANalyzeR [RN55]).
Few papers used a common NLP embedding model other than word2vec in their research. [RN67] used RoBERTa within their process to learn semantic relationships within assembly code and passed these embeddings to a Message Passing Neural Network (MPNN) for vulnerability detection. [gallagherllvm] custom-built a RoBERTa model with LLVM code to learn the code-level semantics to detect vulnerable code. [han2022binary] trained a BERT model with pcode and utilized the generated embeddings to train machine learning models to differentiate between vulnerable and non-vulnerable code. Finally, [koo2021semantic] employed BERT and [10.1145/3460120.3484587] used a GPT model to create embeddings for code similarity detection.
Recurrent neural networks commonly used in natural language processing are regularly used to detect vulnerabilities in compiled binaries. For example, [RN36, RN40, RN10, RN9, RN57, RN47, RN51, RN58, schaad2022deeplearningbased] used an LSTM neural network and [RN9, RN42, RN57, schaad2022deeplearningbased] applied a GRU neural network for vulnerability detection. Furthermore, [RN7, RN10, RN9, RN57, RN45, RN51, RN55, schaad2022deeplearningbased] used a BLSTM, and [RN10, RN9, RN57, RN45, RN46, RN51, RN55, schaad2022deeplearningbased] used a BGRU neural network as part of the research they performed.
The papers most similar to the experiment performed within the current paper are [RN9] and [schaad2022deeplearningbased]. Both papers lift assembly language to LLVM using RetDec, perform preprocessing on LLVM, use word2vec to create embeddings, and use a neural network to train a neural network to discover vulnerabilities in the LLVM code. This paper builds on their research by expanding the study to include additional embedding models (BERT and RoBERTa).
VI Future Research
One of the most significant limitations encountered during the current study was the low number of training samples. Once duplicate functions were removed, the number of functions available to train embedding models and LSTM neural networks decreased drastically. This could indicate that the preprocessing phase may have resulted in many samples with slight code nuances being flattened to a single form. However, this could also suggest that the number of functions that did nothing other than call secondary functions comprised a much larger portion of the dataset than originally expected. Future research will re-evaluate the preprocessing performed on each function to determine how much preprocessing resulted in the sample decrease.
It is postulated that the BERT and RoBERTa embeddings may perform better with more data, so significant emphasis in future research will be on identifying ways to increase the number of quality samples used to train the embedding models. Furthermore, it is possible that additional vulnerable and non-vulnerable code samples could also affect LSTM neural network performance. Thus, a second emphasis for future research would be to identify ways to increase the number of code samples that could be used to train these models.
Future research will increase its focus on unidirectional embedding models, including GPT-1, and GPT-2. A different NLP embedding model may achieve better results when there are fewer samples for training embedding models. Furthermore, some researchers [RN7] were able to achieve accuracy gains when adding bidirectionally to their neural networks. Thus, future research will evaluate different neural networks to identify the best performance. Finally, the current study strictly focused on buffer overflow vulnerabilities. However, it would make sense to expand the research to include additional vulnerability categories in the future.
Finally, the study [arp2022and] states that data snooping is one of the biggest pitfalls in current research on the application of machine learning to cyber security. Data snooping occurs when training and validation data are not properly separated. One may argue that data snooping is present within the current paper because all code samples, including the samples used to train the neural networks, were used to train the embedding models. However, it should be noted that strict data segmentation was used when training each neural network. However, the effects of using these data samples when training the embedding models are unknown and will be explored further in later research.
VII Conclusion
The most significant contribution of the current research is that it is the first study to explore the use of multiple bidirectional transformer (i.e. BERT & RoBERTa) embeddings built using LLVM code to train neural networks to detect vulnerabilities in the lifted code. In addition, it explored the effects of Adam and SGD optimizers, and compared the performance of models using varied values for momentum. This study also identified vital limitations that more complex NLP embeddings introduce to the process of using embeddings for vulnerability discovery in compiled binaries, in contrast to simpler word2vec embeddings. For example, increased embedding max_lengths results in significant memory consumption that limits the maximum number of tokens that can be processed in any code block.
At the end of the project, it was determined that, when using embeddings generated from NLP models trained using a limited number of samples (i.e., 118K), the LSTM built using word2vec: CBOW embeddings with the Adam optimizer and an initial learning rate of 0.0001 achieved the highest accuracy (92.3%). Furthermore, although LSTMs trained using word2vec: Skip-Gram embeddings struggled to achieve accuracy gains with an SGD optimizer and lower learning rates, it achieved a 90.6% accuracy when the Adam optimizer was used with the same initial learning rate. This accuracy was higher than the highest accuracy that was achieved by any LSTM neural network trained using BERT or RoBERTa embeddings. The results suggest that complex contextual NLP embeddings may not provide advantages over simpler word2vec models for this task when these models are trained using smaller datasets. The comparative results provide novel insights into selecting optimal embeddings for learning compiler-independent semantic code representations to advance machine learning detection of vulnerabilities in compiled code.
[…]